# Curse of High Dimensionality Issue in Transformer for Long-context Modeling

### v3

@@ 4.2 Redundancy Analysis from Coding Perspective
One naive approach is to add some penalties during optimization. Unfortunately, each context $C({\bf x})$ has its optimal weight within the whole tokens of the context. To address this, benefiting from the group sparsity (Huang & Zhang, 2010; Lounici et al., 2011), we propose a group coding strategy, reducing the inherent redundancy by partitioning the attention weights $\alpha$ into $k$ groups, i.e., $\bar{\alpha}\in\mathbb{R}^{k}$. Each group shares a common weight, effectively pooling the contributions of grouped tokens. Let $G_{1},\ldots,G_{k}$ denote the index sets from the token indices =${\mathcal{I}}=\{1,\ldots,L\}$, where ${\mathcal{I}}=\bigcup_{g=1}^{k}G_{g}$ and $G_{i}\neq G_{j}$ for $\forall i\neq j$. Formally, we define the group weight as $\bar{\alpha}_{g}=\frac{1}{|G_{g}|}\sum_{i\in G_{g}}\alpha_{i}$. The optimization problem is then reformulated as:
@@
$\min_{\bar{\alpha}}\left\|\sum_{g=1}^{k}\frac{\bar{\alpha}_{g}}{|G_{g}|}\!\sum_{j\in G_{g}}{\bf V}_{j}{-}{\bf y}\right\|_{2}^{2},~{}\mbox{s.t.},\sum_{g=1}^{k}\alpha_{g}{=}1,~{}\alpha_{g}{>}0.$ (10)
@@ 5.1 Grouping for Self-attention
where ${\bf P}_{i,\cdot}{=}\mathrm{softmax}({\bf Q}_{\mathrm{max}G_{i}}{\bf K}^{\top}_{[\mathrm{min}G_{i}:\mathrm{max}G_{i}]}){\in}\mathbb{R}^{m}$ is the weight vector of $i$-th group, $G_{i}$ is the index set of $i$-th group, $\min G_{i}$ and $\max G_{i}$ are the first and last token index in the $i$-th group, $r{=}|{\mathcal{T}}^{\mathrm{foc}}|$ and $k{=}(L-r)/m$ are the number of focal tokens and groups. Here, we compute the weights ${\bf P}$ using only the query of the last token in each group. If non-focal tokens are not divisible by $m$, we will include more tokens in ${\mathcal{T}}^{\mathrm{foc}}$, ensuring all tokens are processed.
Note that some tokens may not access preceding tokens after grouping aggregation. For instance, in the rightmost of Figure 1, the query ${\bf Q}_{2}$ cannot attend to the keys ${\bf K}_{1}$ and ${\bf K}_{2}$ (i.e., their masks are zeros at the $4$-th and $5$-th columns) even if the $2$-th token should access the $1$-th and $2$-th tokens. To address this, we introduce complementary keys ${\bf K}^{c}$ and values ${\bf V}^{c}$ to restore the missing information for the query ${\bf Q}_{i}$ (e.g., adding the keys “${\bf K}_{1}$ and ${\bf K}_{2}$” and the values “${\bf V}_{1}$ and ${\bf V}_{2}$” to the query ${\bf Q}_{2}$ at the $6$-th and $7$-th columns):
@@ 5.2 Fast Focal Token Identification for Grouping
where $\|{\bf A}_{\cdot,i}\|_{0}$ is the number of non-zero elements in the $i$-th column of ${\bf A}$, which will be $L{-}i{+}1$ for the decoder-only based transformer. Similar techniques have been used for the inference of LLMs (Zhang et al., 2023; He et al., 2024).
@@ 6.3 Comparisons on Long-Context Modeling
Results on Longbench-E. We conduct experiments on LongBench-E (Bai et al., 2024), which includes tasks with text lengths up to 31K. Table 1 shows that our DGA-LLM achieves competitive performance across tasks, with an average score of $21.14$ for LLaMA2-7B trained on 8K context-length texts. While our method does not significantly outperform others in average scores, it demonstrates notable advantages in Inter-token Latency (ITL) at 16K context length. Specifically, DGA-LLM achieves an ITL of $28.80$ ms, 2.42×2.42$\times$2.42 × faster than vanilla LLaMA2-7B ($69.70$ ms) and 3.55×3.$5$ faster than LM-Infinite ($102.22$ ms). This highlights the efficiency of our group coding strategy in reducing computational redundancy while maintaining performance, making it particularly suitable for time-sensitive applications. More results on ITL are in Section 6.4.
@@ 6.4 Comparisons on Computational Efficiency
To evaluate the inference efficiency of our DGA, we measure the Inter-token Latency (ITL) of the LLaMA2-7B model during the decoding phase across 4K, 8K, and 16K contexts on LongBench-E using a single A800 GPU. Table 3 shows that DGA significantly reduces inference latency compared to other methods. For instance, at 16K context length, DGA-LLM achieves an ITL of $28.79$ ms, 3.28×3.28$\times$3.28 × faster than MInference ($94.34$ ms) and 2.72×2.72$\times$2.72 × faster than StreamingLLM ($78.28$ ms). Notably, other methods are slower than vanilla self-attention, as they lack optimizations for the generation process. As the context length increases from 4K to 16K, other baseline methods show a sharp ITL rise, indicating a slower generation. In contrast, our DGA-LLM maintains stable and low ITL, demonstrating its efficiency for longer texts (see more detailed complexity analysis in Appendix E). Notably, LongLoRA (Chen et al., 2024) uses $S^{2}$-Attention only during training, reverting to vanilla self-attention in inference, resulting in the same speed. These results highlight the efficiency of DGA-LLM in handling long-context tasks.
@@ 7 Conclusion
In this paper, we address computational redundancy in long-context modeling by first reformulating the sequence modeling as a supervised learning task, enabling a clear understanding of token relevance. Based on this reformulation, we theoretically analyze attention sparsity, showing that only a few tokens significantly contribute to predictions, and propose a group coding strategy to enhance robustness and efficiency. Finally, we propose Dynamic Group Attention (DGA), which uses this strategy to adaptively reduce redundancy while retaining critical token interactions. We empirically validate the effectiveness of DGA on long-context tasks, demonstrating significant reductions in computational costs while maintaining competitive performance.
@@ Acknowledgments
This work was partially supported by the Joint Funds of the National Natural Science Foundation of China (Grant No.U24A20327), Key-Area Research and Development Program Guangdong Province 2018B010107001, and TCL Science and Technology Innovation Fund, China.
@@ A.2 Proof of Theorem 2
$\min_{\bar{\alpha}}\left\|\sum_{g=1}^{k}\!\sum_{j\in G_{g}}\frac{\bar{\alpha}_{g}}{m}{\bf V}_{j}{-}{\bf y}\right\|_{2}^{2},~{}\mbox{s.t.},\sum_{j=1}^{k}\alpha_{j}{=}1,~{}\alpha_{j}{>}0.$ (21)
@@ A.3 Proof of Theorem 3
$\min_{\bar{\alpha}}\left\|\sum_{g=1}^{k}\!\frac{\bar{\alpha}_{g}}{m}\sum_{j\in G_{g}}{\bf V}_{j}{-}{\bf y}\right\|_{2}^{2},~{}\mbox{s.t.},\sum_{j=1}^{k}\alpha_{j}{=}1,~{}\alpha_{j}{>}0.$ (31)
@@ Appendix D Details on Causing Mask in Group-Oriented Attention
[…]
@@ Appendix F Ablation Studies
Effect of group Size $m$. In this experiment, we investigate the impact of group size $m$ on model performance and latency. As shown in Table 4, when varying $m$ from 4 to 64, we observe only a slight degradation in model performance, with perplexity (PPL) increasing from $3.20$ to $3.87$ and accuracy decreasing from $70.17$% to $65.00$%. However, computational efficiency shows a notable improvement, with latency significantly decreasing from $189.80$ ms to $162.02$ ms. Thus, we set $m=16$ to strike a balance between computational efficiency and model performance as the default training setting.
@@
Effect of focal token identification. In this experiment, we investigate the effect of different focal token identification strategies on model performance. From Table 5, compared to the randomly selected strategy, our Top-K approach achieves an improvement in accuracy from $66.91$% to $67.00$%. At the same time, perplexity (PPL) shows a decrease from $3.60$ to $3.59$. These results suggest the effectiveness of focal token identification in DGA over the randomly selected strategy.
@@
Effect of importance rate $\gamma$. In this experiment, we investigate the effect of the importance rate $\gamma$ in our DGA. Table 6 demonstrates a significant positive correlation between $\gamma$ values and model performance. Specifically, as $\gamma$ increases from 0.1 to 0.9, the model accuracy improves substantially from $67.00$% to $70.58$%, while perplexity (PPL) consistently decreases from $3.59$ to $3.20$. However, the latency shows an increase as $\gamma$ increases, growing from $172.12$ ms to $555.25$ ms. A larger $\gamma$ indicates that more tokens are identified as focal tokens, thereby preserving a greater portion of the original input information in the self-attention computations. This preservation of information, while beneficial for accuracy, inevitably increases the computational overhead. Conversely, a smaller $\gamma$ leads to more tokens being grouped and aggregated, which enhances computational efficiency but at the expense of model performance. Striking a balance between computational efficiency and model effectiveness, we select $\gamma=0.1$ as the default training setting.
@@
Effect of complementary tokens. Table 7 shows that removing complementary tokens significantly degrades performance on tasks like Multi-Doc QA from 3.58 to 2.37 and Code from $53.45$ to $48.00$, indicating complementary tokens are critical for context modeling. They slightly increase latency from $24.9$ ms to $28.8$ ms, but still 2.43.5×2.4$\sim$ 3.5$\times$2.4 ∼ 3.5 × faster than vanilla self-attention (Table 2: $69.70\sim 102.22$ ms).

### v4

@@ 4.2 Redundancy Analysis from Coding Perspective
One naive approach is to add some penalties during optimization. Unfortunately, each context $C({\bf x})$ has its optimal weight within the whole tokens of the context. To address this, benefiting from the group sparsity (Huang & Zhang, 2010; Lounici et al., 2011), we propose a group coding strategy, reducing the inherent redundancy by partitioning the attention weights $\alpha$ into $k$ groups, i.e., $\bar{\alpha}\in\mathbb{R}^{k}$. Each group shares a common weight, effectively pooling the contributions of grouped tokens. Let $G_{1},\ldots,G_{k}$ denote the index sets from the token indices ${\mathcal{I}}=\{1,\ldots,L\}$, where ${\mathcal{I}}=\bigcup_{g=1}^{k}G_{g}$ and $G_{i}\bigcap G_{j}=\emptyset$ for $\forall i\neq j$. Formally, we define the group weight as $\bar{\alpha}_{g}=\frac{1}{|G_{g}|}\sum_{i\in G_{g}}\alpha_{i}$. The optimization problem is then reformulated as:
@@
$\min_{\bar{\alpha}}\left\|\sum_{g=1}^{k}\frac{\bar{\alpha}_{g}}{|G_{g}|}\!\sum_{j\in G_{g}}{\bf V}_{j}{-}{\bf y}\right\|_{2}^{2},~\mbox{s.t.},\sum_{g=1}^{k}\bar{\alpha}_{g}|G_{g}|{=}1,~\alpha_{g}{>}0.$ (10)
@@ 5.1 Grouping for Self-attention
where ${\bf P}_{i,\cdot}{=}\mathrm{softmax}({\bf Q}_{\mathrm{max}G_{i}}{\bf K}^{\top}_{[\mathrm{min}G_{i}:\mathrm{max}G_{i}]}){\in}\mathbb{R}^{m}$ is the weight vector of $i$-th group, $G_{i}$ is the index set of $i$-th group, $\min G_{i}$ and $\max G_{i}$ are the first and last token index in the $i$-th group, $r{=}|{\mathcal{T}}^{\mathrm{foc}}|$ and k=(Lr)/mk{=}(L-r)/m are the number of focal tokens and groups. Here, we compute the weights $\bf P$} using only the query of the last token in each group. If non-focal tokens are not divisible by $m$, we will include more tokens in ${\mathcal{T}}^{\mathrm{foc}}$, ensuring all tokens are processed.
Note that some tokens may not access preceding tokens after grouping aggregation. For instance, in the rightmost of Figure 1, the query ${\bf Q}_{2}$ cannot attend to the keys ${\bf K}_{1}$ and ${\bf K}_{2}$ (i.e., their masks are zeros at the $4$-th and $5$-th columns) even if the $2$-th token should access the $1$-th and $2$-th tokens. To address this, we introduce complementary keys ${\bf K}^{c}$ and values ${\bf V}^{c}$ to restore the missing information for the query ${\bf Q}_{i}$ (e.g., adding the keys ${\bf K}_{1}$ and ${\bf K}_{2}$” and the values ${\bf V}_{1}$ and ${\bf V}_{2}$” to the query ${\bf Q}_{2}$ at the $6$-th and $7$-th columns):
@@ 5.2 Fast Focal Token Identification for Grouping
where $\|{\bf A}_{\cdot,i}\|_{0}$ is the number of non-zero elements in the $i$-th column of $\bf A$}, which will be Li+1L{-}i{+}1 for the decoder-only based transformer. Similar techniques have been used for the inference of LLMs (Zhang et al., 2023; He et al., 2024).
@@ 6.3 Comparisons on Long-Context Modeling
Results on Longbench-E. We conduct experiments on LongBench-E (Bai et al., 2024), which includes tasks with text lengths up to 31K. Table 1 shows that our DGA-LLM achieves competitive performance across tasks, with an average score of $21.14$ for LLaMA2-7B trained on 8K context-length texts. While our method does not significantly outperform others in average scores, it demonstrates notable advantages in Inter-token Latency (ITL) at 16K context length. Specifically, DGA-LLM achieves an ITL of $28.80$ ms, $2.42\times$ faster than vanilla LLaMA2-7B ($69.70$ ms) and $3.55\times$ faster than LM-Infinite ($102.22$ ms). This highlights the efficiency of our group coding strategy in reducing computational redundancy while maintaining performance, making it particularly suitable for time-sensitive applications. More results on ITL are in Section 6.4.
@@ 6.4 Comparisons on Computational Efficiency
To evaluate the inference efficiency of our DGA, we measure the Inter-token Latency (ITL) of the LLaMA2-7B model during the decoding phase across 4K, 8K, and 16K contexts on LongBench-E using a single A800 GPU. Table 3 shows that DGA significantly reduces inference latency compared to other methods. For instance, at 16K context length, DGA-LLM achieves an ITL of $28.79$ ms, $3.28\times$ faster than MInference ($94.34$ ms) and $2.72\times$ faster than StreamingLLM ($78.28$ ms). Notably, other methods are slower than vanilla self-attention, as they lack optimizations for the generation process. As the context length increases from 4K to 16K, other baseline methods show a sharp ITL rise, indicating a slower generation. In contrast, our DGA-LLM maintains stable and low ITL, demonstrating its efficiency for longer texts (see more detailed complexity analysis in Appendix E). Notably, LongLoRA (Chen et al., 2024) uses $S^{2}$-Attention only during training, reverting to vanilla self-attention in inference, resulting in the same speed. These results highlight the efficiency of DGA-LLM in handling long-context tasks.
@@ 7 Conclusion
In this paper, we address computational redundancy in long-context modeling by first reformulating the sequence modeling as a supervised learning task, enabling a clear understanding of token relevance. Based on this reformulation, we theoretically analyze attention sparsity, showing that only a few tokens significantly contribute to predictions, and propose a group coding strategy. Last, we propose Dynamic Group Attention (DGA), which uses this strategy to adaptively reduce redundancy while retaining critical token interactions. We validate the effectiveness of DGA on long-context tasks, showing significant reductions in computational costs while maintaining competitive performance.
@@ Acknowledgments
This work was partially supported by the Joint Funds of the National Natural Science Foundation of China (Grant No.U24A20327), Key-Area Research and Development Program Guangdong Province 2018B010107001, Postdoctoral Fellowship Program of CPSF (GZC20251043), and TCL Science and Technology Innovation Fund, China.
@@ A.2 Proof of Theorem 2
$\min_{\bar{\alpha}}\left\|\sum_{g=1}^{k}\!\sum_{j\in G_{g}}\frac{\bar{\alpha}_{g}}{m}{\bf V}_{j}{-}{\bf y}\right\|_{2}^{2},~\mbox{s.t.},\sum_{g=1}^{k}m\alpha_{g}{=}1,~\alpha_{g}{>}0.$ (21)
@@ A.3 Proof of Theorem 3
$\min_{\bar{\alpha}}\left\|\sum_{g=1}^{k}\!\sum_{j\in G_{g}}\frac{\bar{\alpha}_{g}}{m}{\bf V}_{j}{-}{\bf y}\right\|_{2}^{2},~\mbox{s.t.},\sum_{g=1}^{k}m\alpha_{g}{=}1,~\alpha_{g}{>}0.$ (31)
@@ Appendix D Details on Causing Mask in Group-Oriented Attention
5: Final mask is ${\bf M}=[{\bf M}^{0}_{{\mathcal{T}}^{\mathrm{foc}}},{\bf M}^{n},{\bf M}^{c}]\in\mathbb{R}^{L\times(r+k+m)}$.
@@ Appendix F Ablation Studies
Effect of group Size $m$. In this experiment, we investigate the impact of group size $m$ on model performance and latency. As shown in Table 4, when varying $m$ from 4 to 64, we observe only a slight degradation in model performance, with perplexity (PPL) increasing from $3.20$ to $3.87$ and accuracy decreasing from 70.1770.17% to 65.0065.00%. However, computational efficiency shows a notable improvement, with latency significantly decreasing from $189.80$ ms to $162.02$ ms. Thus, we set $m=16$ to strike a balance between computational efficiency and model performance as the default training setting.
@@
Effect of focal token identification. In this experiment, we investigate the effect of different focal token identification strategies on model performance. From Table 5, compared to the randomly selected strategy, our Top-K approach achieves an improvement in accuracy from 66.9166.91% to 67.0067.00%. At the same time, perplexity (PPL) shows a decrease from $3.60$ to $3.59$. These results suggest the effectiveness of focal token identification in DGA over the randomly selected strategy.
@@
Effect of importance rate $\gamma$. In this experiment, we investigate the effect of the importance rate $\gamma$ in our DGA. Table 6 demonstrates a significant positive correlation between $\gamma$ values and model performance. Specifically, as $\gamma$ increases from 0.1 to 0.9, the model accuracy improves substantially from 67.0067.00% to 70.5870.58%, while perplexity (PPL) consistently decreases from $3.59$ to $3.20$. However, the latency shows an increase as $\gamma$ increases, growing from $172.12$ ms to $555.25$ ms. A larger $\gamma$ indicates that more tokens are identified as focal tokens, thereby preserving a greater portion of the original input information in the self-attention computations. This preservation of information, while beneficial for accuracy, inevitably increases the computational overhead. Conversely, a smaller $\gamma$ leads to more tokens being grouped and aggregated, which enhances computational efficiency but at the expense of model performance. Striking a balance between computational efficiency and model effectiveness, we select $\gamma=0.1$ as the default training setting.
@@
Effect of complementary tokens. Table 7 shows that removing complementary tokens significantly degrades performance on tasks like Multi-Doc QA from 3.58 to 2.37 and Code from $53.45$ to $48.00$, indicating complementary tokens are critical for context modeling. They slightly increase latency from $24.9$ ms to $28.8$ ms, but still $2.4\sim 3.5\times$ faster than vanilla self-attention (Table 2: $69.70\sim 102.22$ ms).